1. Given a firm’s demand function, P = 24 - 0.5Q and the average cost function, AC = Q 2 – 8Q + 36 + 3/Q, calculate the level of output Q which a) maximizes total revenue b) maximizes profits


1.     Given a firm’s demand function, P = 24 - 0.5Q and the average cost function, AC = Q2 – 8Q + 36  + 3/Q, calculate the level of output Q which  a)  maximizes total revenue


       b) maximizes profits


Grade:

1 Answers

Priyansh Bajaj AskiitiansExpert-IITD
13 years ago

Dear Sanjay,

Solution:- (a) Total Revenue (TR) = P*Q = (24 - 0.5Q) * Q

To find the max value of Total Revenue, differentiate it and equate it to zero.

d(TR)/dQ = 0  or  d(24Q - 0.5Q2)/dQ = 0

24 - Q = 0 or Q = 24 (ANS)

(b) Total cost (TC) = (AC) * Q = (Q2 - 8Q +36 +3/Q) * Q = (Q3 - 8Q2 +36Q +3)

Profit (P) =  [TR - TC] = [(24Q - 0.5Q2) - (Q3 - 8Q2 +36Q +3)] = (-12Q +7.5Q2 -Q3 -3)

To find the max value of Profit, differentiate it and equate it to zero.

d(P)/dQ = 0  or  d(-12Q +7.5Q2 -Q3 -3)/dQ = 0

-12 +15Q -3Q2 = 0 or Q2- 5Q+4 =0

Q =1 or 4

At Q =1, P = - 8.5 and at Q = 4, P = 5

So, 'P' is maximum at Q =4 (ANS)



Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best Sanjay!!!


Regards,

Askiitians Experts

Priyansh Bajaj
?